1 2 77 (9 pts) 7. Find 4, by hand, if A = 1 2 5 1-1 1 2
Answer #1

entry at row 2, Make ceros in column 2 except the columa (pivot entry) Ra o 27 13 0-2 .A o 0 RI → Ri -2R2 0 O 1 0 1 3 - o 7.4solution A 다. 1 2 1413 5 2 7 5 -1 12 - 2 = 1(4-5) -2 (2-(-5)) ++(1-1-2)) =!(-1) - 2 (2+5) +7(1+2) =\(-1)-267) + +(3) -1-14 +2